OpenStudy (anonymous):

Part a.) Write the equation of the line that has a slope of -2 and a y-intercept of -6. (5 points) Part b.) Write the equation of the line that passes through the point (6, -2) and has a slope of -2/3. (5 points)

OpenStudy (4everaddicted2anime):

For part a the formula is y=mx+b Plug in the slope for m and the y-intercept for b to get the answer

OpenStudy (4everaddicted2anime):

For part b you need to plug in what the question gives you into the formula y=mx+b \[-2= \frac{ -2 }{ 3 }*6+b\] Now solve for b